OpenStudy (cutepuppies):

Celia lives 2/3 mile from the train station. Dino lives 1/5 the distance that Celia lives from the station. How far does Dino live from the train station? miles

OpenStudy (jdoe0001):

1/5 of 2/3 -> 2/3 * 1/5

OpenStudy (anonymous):

2/15

OpenStudy (qwertty123):

Step 1.) 1/5 of 2/3....." of " means multiply 1/5 * 2/3 =?

OpenStudy (iluvsoccer):

\[\frac{ 1 }{ 5 }\times \frac{ 2 }{ 3 }\]

OpenStudy (iluvsoccer):

Just simply multiply straight across. :)

OpenStudy (cutepuppies):

2/15

OpenStudy (jdoe0001):

and so @cutepuppies what did you get anyway?

OpenStudy (iluvsoccer):

Good. :)

OpenStudy (jdoe0001):

2/15 \(\huge \checkmark\)

OpenStudy (qwertty123):

YEAH! :D

OpenStudy (cutepuppies):

is it 36/4

OpenStudy (cutepuppies):

it has to be a hole number